OpenStudy (anonymous):

Multiply and simplify the product [2sqrt(7) - sqrt(3)]2

OpenStudy (anonymous):

4root(7)-2root(3)

OpenStudy (anonymous):

31 - 4sqrt21

OpenStudy (anonymous):

\[2(2\sqrt{7}-\sqrt{3})=4\sqrt{7}-2\sqrt3\]